Follow the steps below to estimate 1 with an error less than TOO, Let f(z) = ln(1+2). In class on March 15, we showed that fo please show the steps! than you!
Answer #1

Let f(z) = ln(1+1). for all n 21, f(n)(z) =ー(1+2)n (-1)-1(n-1 )! (1) For each n, find the maximum value Ma+l of tna) for zin